What is the total surface area of the walls of a hospital ward 5m long, 4m wide and 2.5m high.

Answer by math_tutor2020(3817)


Answer: 45 square meters

Work Shown:
L = 5 = length
W = 4 = width
H = 2.5 = height
SA = total surface area of all walls + floor + ceiling
SA = 2*(LW + LH + WH)
SA = 2*(5*4 + 5*2.5 + 4*2.5)
SA = 85
LSA = lateral surface area = area of the walls only (i.e exclude the floor and ceiling)
LSA = SA - (floor area) - (ceiling area)
LSA = 85 - (5*4) - (5*4)
LSA = 85 - 20 - 20
LSA = 45 square meters is the area of all the walls only.

Another approach:
The perimeter of the rectangular floor is 2*(L+W) = 2*(5+4) = 18 meters. Multiply this with the height to get 18*2.5 = 45 square meters

Yet another approach:
The 5 by 2.5 walls are 5*2.5 = 12.5 sq meters each. So far that accounts for 12.5+12.5 = 25 square meters of wall area.
The 4 by 2.5 walls have area of 4*2.5 = 10 square meters each, subtotaling 10+10 = 20 additional square meters of wall.
The total wall area is 25+20 = 45 square meters